OpenStudy (anonymous):

Solve the system of equations. 5x + y = 9 3x + 2y = 4 (-2, 5) (1, 4) (2, -1) (4, -4) Determine which system below will produce infinitely many solutions. 2x + 5y = 24 2x + 5y = 42 3x - 2y = 15 6x + 5y = 11 4x - 3y = 9 -8x + 6y = -18 5x - 3y = 16 -2x + 3y = -7 Choose the equivalent system of linear equations that will produce the same solution as the one given below. 6x + 2y = -6 3x - 4y = -18 12x + 4y = -12 15x = -30 8x + 4y = -4 14x = -10 6x - 8y = -36 -6y = -42 6x - y = -15 3y = 9

OpenStudy (anonymous):

need help

OpenStudy (igreen):

Okay, no problem.

OpenStudy (anonymous):

thanx you

OpenStudy (igreen):

Solve the system of equations. \(5x + y = 9\) \(3x + 2y = 4\) Solve the first equation for \(y\). \(5x + y = 9\) Subtract \(-5x\) to both sides: \(y = -5x + 9\) Now we can input \(-5x + 9\) for \(y\) in the 2nd equation: \(3x + 2y = 4\) \(3x + 2(-5x+9) = 4\) Distribute \(2\) into the parenthesis: \(3x - 10x + 18 = 4\) Simplify: \(-7x + 18 = 4\) Subtract \(18\) to both sides: \(-7x = 4 - 18\) Simplify: \(-7x = -14\) Divide \(-7\) to both sides: \(x = 2\) Okay, now we have our x-value to the solution. We can re-plug this into any of the two equations to find the y-value to the solution: \(5x + y = 9\) \(5(2) + y = 9\) \(10 + y = 9\) Subtract \(10\) to both sides: \(y = 9 - 10\) SImplify: \(y = -1\) So our solution is \((2, -1)\).

OpenStudy (igreen):

@alenahall12

OpenStudy (anonymous):

i review the work you did

OpenStudy (igreen):

\(\bf Determine~which~system~below~will~produce~infinitely~many~solutions.\) When a solution of two lines produce infinitely many solutions those two lines will be the same line.

OpenStudy (igreen):

So..look at the answers and find out which one is the same line. \(\bf Option~A \! \! \! :\) \(2x + 5y = 24\) \(2x + 5y = 42\) These last two numbers are switched around. You can tell these lines are parallel. \(\bf Option~B \! \! \! :\) \(3x - 2y = 15\) \(6x + 5y = 11\) These obviously have one solution..they are clearly different. \(\bf Option~C \! \! \! :\) \(4x - 3y = 9\) \(-8x + 6y = -18\) This is the correct answer! These lines are the same because the 2nd equation is the first equation multiplied by 2.

OpenStudy (anonymous):

hum i see that makes some sense

OpenStudy (igreen):

\(\bf Choose~the~equivalent~system~of~linear~equations~that~will~produce~the~same\) \(\bf solution~as~the~one~given~below.\) \(6x + 2y = -6\) \(3x - 4y = -18\) I'd just use a graph for this..it will take too much time to just find out the solution then check the others. I recommend this one: https://www.desmos.com/calculator Just plug the equations in separate boxes on the left. In this case, the solution is \((-2, 3)\). Option A's solution is also \((-2,3)\).

OpenStudy (igreen):

So your answer is A. @alenahall12

OpenStudy (anonymous):

oh sorry i was writing this'll down in notes it really helps when u go in the steps

OpenStudy (anonymous):

and thank you big help

OpenStudy (igreen):

No problem.